OpenStudy (anonymous):

Let S be the portion of the paraboloid z = 4 -x^2 -y^2 that lies above z= 0. Let F be the vector field F = (z-y)i + (x+z)j - (e^xyz(cosy))k. Using Stoke's Theorem, find the surface integral.

OpenStudy (anonymous):

So the hope is that \(\nabla \times \mathbf F\) is a very simple function, if we're using Stokes.

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

Can you calculate it?

OpenStudy (anonymous):

Yes. Let me re-calculate to make sure my answer isn't fudged.

OpenStudy (anonymous):

\[(xze ^{xyz}cosy -e^{xyz}siny - 1)i -(e^{xyz}cosy - 1)j + 2k\]

OpenStudy (anonymous):

What a mess.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay, do you know how to parametrize the parabaloid then?

OpenStudy (anonymous):

Maybe...

OpenStudy (anonymous):

subtract z in the equation, and take the gradient of the expression.

OpenStudy (anonymous):

-2x-2y, or -2x-2y-1? I dont quite understand what you meant by subtract z or why it's done that way.

OpenStudy (anonymous):

the gradient is a vector

OpenStudy (anonymous):

When you have a level set \[ g(x,y,z) = C \]Then the gradient will give you the normal vector to that level set.

OpenStudy (anonymous):

So you're saying that \[g(x,y,z) = C = 4 - x^2 -y^2 - z\] \[gradient = (-2x)i - (2y)j - 1k \] ? Sorry, this is again unfamiliar to me.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

you could have even said \[ g(x,y,z)=-4 = -x^2-y^2-z \]It doesn't matter since \(C\) goes away when differentiated.

OpenStudy (anonymous):

Excellent

OpenStudy (anonymous):

Seems like a tough integral...

OpenStudy (anonymous):

Would you use \[\int\limits_{}^{}\int\limits_{S}^{} (curlF) ndS = \int\limits_{C}^{} Fdr \]

OpenStudy (anonymous):

Where \[Fdr = (z-y)dx + (x+z)dy + (e^{xyz}cosy)dz\]

OpenStudy (anonymous):

In doing so, you are no longer using stokes.

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

\[dS = \sqrt(4x^2 + 4y^2 +1) ??\]

OpenStudy (anonymous):

No, okay, so basically \(\mathbf n\) is the unit normal vector, and \(dS\) is the magnitude of the the normal vector \(d\mathbf S\), which you just calculated.

OpenStudy (anonymous):

Well actually, I'm being a bit sloppy here.

OpenStudy (anonymous):

Anyway \(d\mathbf S = \langle -2x,-2y,-1\rangle dxdy\)

OpenStudy (anonymous):

And \[ \mathbf n~ dS = d\mathbf S \]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So do we leave the z's as constants the whole time? or set all z's to 0?

OpenStudy (anonymous):

Okay, so basically, the \(z\) gets replaced with \(z(x,y)\) when we do this sort of parametrization.

OpenStudy (anonymous):

Since \(z=f(x,y)\).

OpenStudy (anonymous):

So, for every z, we replace with 4 -x^2 - y^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you sure you wrote the problem down correctly?

OpenStudy (anonymous):

Very sure.

OpenStudy (anonymous):

Our only hope is that cylindrical coords make things easier.

OpenStudy (anonymous):

\[(x(4-x^2-y^2)e^{((4-y^2-x^2)xy)}cosy- e^{(4-x^2-y^2)yx}siny -1 )i -(y(4-x^2-y^2)e^{(4-x^2-y^2)yx}cosy-1)j +2k\]

OpenStudy (anonymous):

Doesnt even fit. -_-

OpenStudy (anonymous):

How should I go about using cyl coords for this? The only I've been to taught to use this is to convert it to a line integral since Stokes' Theorem allows for you to choose whichever is more convenient, which seems like the line integral would be for this particular instance

OpenStudy (anonymous):

I should have said polar coords. \[ r^2 = x^2+y^2\\ \theta = \tan\frac yx\\ x = r\cos\theta\\ y = r\sin\theta \\ dx~dy = r~dr~d\theta \]

OpenStudy (anonymous):

Bleh. Either conversion is just... tedious. But given that....\[(\cos(\theta)(4r-r^3) e^{(4r^2-r^4)\cos(\theta)\sin(\theta)}\cos(rsin(\theta)) - e^{(4r^2-r^4)\cos(\theta)\sin(\theta)}\sin(rsin(\theta) - 1)i -(\sin(\theta)(4r-r^3)e^{(4r^2-r^4)\cos(\theta)\sin(\theta)}-1)j +2k\]

OpenStudy (anonymous):

There is something wrong with this problem.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I swear that F(x,y,z) = (z-y)i +(x+z)j - (e^xyz (cosy))k. S is the portion of the paraboloid z = 4 - x^2 - y^2 that lies above z=0.

OpenStudy (anonymous):

Given Stokes' Theorem, why can't you use a line integral? Stokes' Theorem just equates them. Using it either way is using the theorem. Yes?

OpenStudy (anonymous):

http://www.math.ucsd.edu/programs/undergraduate/math_20E_exam/Practice_Test.pdf would be the link to the question also. It's the 3rd question.

OpenStudy (anonymous):